OpenStudy (anonymous):

is 8:12,12:18 a proportion

OpenStudy (anonymous):

8/12 = 2/3 (divided all by 4) 12/18 = 2/3 (divided all by 6) proportion!!

OpenStudy (anonymous):

@midnightwolf good luck :)